[24,3,9] , [38,8,11] , [40,13,7] , [?,1,3]
What is the missing number?

Answer by ikleyn(52787)

.
The pattern is obvious:


[24,3,9]  :  24 = 2*(3+9)

[38,8,11] :  38 = 2*(8+11)

[40,13,7] :  40 = 2*(13+7)

[?,1,3]   :   ? = 2*(1+3) = 8.